# Harmonic analysis of atomic vapor magnetization for in situ calibration of tri-axial time-dependent magnetic field

### v2

@@ IV.1 Static field alignment
We have then applied an additional transverse static component that shifts the free precession frequency to 4950 Hz: this corresponds to have $(m_{1},m_{2},m_{3})\approx(0,52,66)\cdot 10^{-3}$, which according to eq.(4) causes a second-harmonic peak about 34 dB below the fundamental tone. This expectation is in perfect accordance with the spectrum shown in Fig.5b. For comparison, the condition of Fig.4, with the second-harmonic peak 50 dB weaker than the fundamental tone (hence $A_{2}\approx 10^{-5}$) gives $m_{1}^{2}+m_{2}^{2}\approx 8\cdot A_{2}\approx 8\cdot 10^{-5}$. Despite this small value, as $m_{3}$ is also small, this corresponds to a $8\degree$ misalignment. This value would decrease down to $2.6\degree$ in case of a 60 dB ratio, i.e. if the second-harmonic peak were at the level of the noise floor.
Figure 5: As in Fig.4, the magnetization adiabatically follows a 16 $\mu$T, 3.144 kHz field rotating on the $xy$ plane. But a transverse component (830 nT in amplitude) is added to the static field along the $z$ direction. Also in this case the signal is plotted in the time (a) and frequency (b) domains. Compared to Fig.4b a second-harmonic peak emerges, in quantitative accordance with the prediction of eq.(4). The red, blue, orange, and green lines indicate the field rotation frequency and its $2^{\mathrm{nd}}$, $3^{\mathrm{rd}}$, and $4^{\mathrm{th}}$ harmonics, respectively.
@@
IV.2 Effects of polarization ellipticity
@@
More generally, the method can be applied to identify the phase difference between the drive signal and the magnetic field generated by each coil making it possible to generate fields with an arbitrary configuration.
V Conclusion
@@
We have shown that second harmonic terms can be used to point out the presence of DC field components lying on the polarization plane of the rotating one. Then we have described a procedure suited to refine the circular polarization of the latter. In typical conditions, the first goal can also be pursued with a traditional approach based on Larmor frequency minimization. In contrast, the method developed for polarization refinement constitutes a useful tool to retrieve accurate calibration factors, which enable a precise set of relative amplitudes and phases of fast-oscillating field components.
We have developed and examined the procedure to produce a circularly polarized field on a given plane. Analogous procedures can be implemented on a perpendicular plane and at the diverse frequencies of interest. This would provide a complete set of calibration factors sufficient to generate three-dimensional fields with arbitrary configuration and time dependence.
@@
As discussed in Sec.III, the proposed methodology (developed to point out spurious components of the static field and polarization imperfections of the time-dependent one) is based on the harmonic analysis of the detected signal, which is proportional to the $x$ component of $\hat{u}_{B}(t)=\vec{B}(t)/B(t)$ that is
$\begin{split}S(t)&=(m_{1}+\cos(\omega t))\big{[}\big{(}m_{1}+\cos(\omega t)\big{)}^{2}+\\ &+\big{(}m_{2}+(1+\varepsilon)\sin(\omega t+\phi)\big{)}^{2}+m_{3}^{2}\big{]}^{-1/2}\end{split}$ (A.1)
A multivariate Taylor expansion of $S(t)$ in the small variables $m_{1},m_{2},\epsilon$ and $\phi$ followed by a Fourier analysis leads to determining the signal components at $\omega$ and its multiples.

### v3

@@ 4.1 Static field alignment
We have then applied an additional transverse static component that shifts the free precession frequency to 4950 Hz: this corresponds to have $(m_{1},m_{2},m_{3})\approx(0,52,66)\cdot 10^{-3}$, which according to eq.4 causes a second-harmonic peak about 34 dB below the fundamental tone. This expectation is in perfect accordance with the spectrum shown in Fig.5b. For comparison, the condition of Fig.4, with the second-harmonic peak 50 dB weaker than the fundamental tone (hence $A_{2}\approx 10^{-5}$) gives $m_{1}^{2}+m_{2}^{2}\approx 8\cdot A_{2}\approx 8\cdot 10^{-5}$, this corresponds to a $8\degree$ misalignment, which would decrease down to $2.6\degree$ in case of a 60 dB ratio, i.e. if the second-harmonic peak were at the level of the noise floor. It is worth noting that the angular uncertainty is not a reliable measure of the method’s accuracy because, for small values of $m_{3}$ (as in this case), even minor uncertainties in $m_{1},m_{2}$ result in significant angular misalignments.
Figure 5: As in Fig.4, the magnetization adiabatically follows a 16 $\mu$T, 3.144 kHz field rotating on the $xy$ plane. But a transverse component (830 nT in amplitude) is added to the static field along the $z$ direction. Also in this case the signal is plotted in the time (a) and frequency (b) domains. Compared to Fig.4b a second-harmonic peak emerges, in quantitative accordance with the prediction of eq.4. The red, blue, orange, and green lines indicate the field rotation frequency and its $2^{\mathrm{nd}}$, $3^{\mathrm{rd}}$, and $4^{\mathrm{th}}$ harmonics, respectively.
@@
4.2 Effects of polarization ellipticity
@@
More generally, the method can be applied to identify the phase difference between the drive signals and the magnetic field generated by the corresponding coils, making it possible to produce arbitrary 3D magnetic fields.
5 Conclusion
@@
We have shown that second harmonic terms can be used to reveal the presence of DC field components lying on the polarization plane of the rotating one. Then we have described a procedure suited to refine the polarization of the latter. In typical conditions, the first goal can also be pursued with a traditional approach based on Larmor frequency minimization. In contrast, the method developed for polarization refinement constitutes a useful tool to retrieve accurate calibration factors taking into account spurious effects as parasitic capacitance or eddy currents induced in the surroundings, which enable the application of fast-oscillating field components with precisely assigned relative amplitudes and phases.
We have examined the procedure for the specific case of producing a circularly polarized field on a given plane at a given frequency. An analogous procedure can be implemented on a perpendicular polarization plane and at the diverse frequencies of interest. This would provide a complete set of calibration factors sufficient to generate three-dimensional fields with arbitrary configuration and time dependence.
Acknowledgments
A.Fregosi is pleased to acknowledge the support of EuPRAXIA Advanced Photon Sources project under the contract EuAPS IR0000030, CUP I93C21000160006. G.Bevilacqua acknowledges the partial support of the MIUR Project PRIN 2020, Mathematics for Industry 4.0, Project No. 2020F3NCPX, and of the GNFM-INdAM.
@@ Acknowledgments
As discussed in Sec.3, the proposed methodology (developed to point out spurious components of the static field and polarization imperfections of the time-dependent one) is based on the harmonic analysis of the detected signal, which is proportional to the $x$ component of $\hat{u}_{B}(t)=\vec{B}(t)/B(t)$ that is
$S(t)=\frac{m_{1}+\cos(\omega t)}{\sqrt{[m_{1}+\cos(\omega t)]^{2}+[m_{2}+(1+\varepsilon)\sin(\omega t+\phi)]^{2}+m_{3}^{2}}}$ (A.1)
A multivariate Taylor expansion of $S(t)$ in the small variables $m_{1},m_{2},\varepsilon$ and $\phi$ followed by a Fourier analysis leads to determining the signal components at $\omega$ and its multiples.